# Reorganizing attention-space geometry with expressive attention

### v2

@@ Abstract
Attention regulates information transfer between tokens. For this, query and key vectors are compared, typically in terms of a scalar product, $\mathbf{Q}^{T}\mathbf{K}$, together with a subsequent softmax normalization. In geometric terms, the standard dot-product attention (DPA) leads to large/small attention weights for parallel/antiparallel queries and keys. Here we study expressive attention (EA), which is based on $(\mathbf{Q}^{T}\mathbf{K})^{2}$, the squared dot product. In this case, attention is enhanced when query and key are either parallel or antiparallel, and suppressed for orthogonal configurations. EA can be introduced into any attention-based code without additional compute costs or memory requirements. For a series of autoregressive prediction tasks, we find that expressive attention performs at least as well as vanilla DPA. Increasing task complexity, EA is observed to outperform DPA with increasing margins, which also holds for multi-task settings. For a given model size, EA manages to achieve 100% performance for a range of complexity levels not accessible to DPA. Our results show that it is possible to reorganize the geometry of the matching condition in the space of attention heads without loss of performance.
1 Introduction
@@
Since its inception (?, ?), the attention mechanism has had a transformative impact on machine learning (?, ?). At its core, attention facilitates pairwise information routing between tokens, with information being transmitted when a given matching condition is fulfilled. For the latter, query and key vectors $\mathbf{Q}$ and $\mathbf{K}$ are compared in terms of the respective dot products, $\mathbf{Q}^{T}\mathbf{K}$. In the standard formulation, denoted here dot-product attention (DPA), a softmax operation is used to map the scalar products between queries and keys to a normalized probability distribution. As illustrated in Fig. 1, dot-product attention constrains the matching condition essentially to a one-dimensional subspace of the otherwise potentially large space of attention heads. Given that attention matrices tend to be sparse (?), it may however be favorable to allow the system to express low attention states in a larger subspace.
@@
Here we introduce and discuss a modified attention mechanism, expressive attention (EA), which is based on the presumption that low attention should correspond to orthogonal query and key configurations. This venue allows attention to express itself in the entire attention space, as illustrated in Fig. 1. Interestingly, EA is realized when using $(\mathbf{Q}^{T}\mathbf{K})^{2}$ for calculating attention weights, viz the squared scalar product. For applications, this approach can be implemented with minimal efforts, normally just a single line of code needs to be modified. Computational complexity and memory requirements are not affected.
@@
We present a comparative evaluation of vanilla DPA and EA. For this purpose we employ a suite of autoregressive prediction tasks (NT prediction tasks) that are based on time series generated by the delayed addition of two or more numbers, where addition is modulo a generic basis $N$, like 2 or 16 (binary/hexadecimal case). NT tasks are generalized delayed XOR tasks that become increasingly difficult to solve for larger $N$ and delays $\tau$ (?). Models are expected to achieve not just a good performance, but 100%, which reflects the view that systems can be said to exhibit reasoning errors when achieving less than 100% accuracy for solvable tasks (?). The question is then, how long perfect performance can be maintained when systematically increasing task difficulty, viz complexity.
@@ 2 Related work
Alternatives to the standard, attention-based transformer have been studied intensively during the past years. At a basic level, non-transformer based architectures for sequence processing have been proposed, like structured and/or selective state sequence models (?, ?). Other approaches aim to reduce compute costs, for example by making use of a multi-to-multi recurrent network analogy (?). Another route is linearized attention (?, ?, ?) for which compute scales linearly with context length, and not quadratically. In practice, linear attention comes with various problems of its own, like unbounded gradients and attention dilution (?, ?). In general, linearized attention models substitute the softmax operation of classical dot-product attention by various linear kernel functions, which may have stochastic components (?, ?), well designed spatial dependencies (?), or being derived from a singular value decomposition (?).
@@
A different aspect is addressed by rotary positional embedding (RoPE), which adds positional information directly to the attention mechanism, via position-dependent rotations of query and key vectors (?, ?). RoPE can be used in particular for extending context windows (?). Somewhat similar in spirit is $(IA)^{3}$, which uses element-wise rescaled key and values vectors (in addition to rescaled hidden layer activities) for fine-tuning pretrained LLMs for downstream tasks (?).
@@
Two main routes are available for testing the performance of sequence processing architectures, such as transformers. The first is to use databases relevant for real-world applications (?, ?), the second is to rely on synthetic test suites. The latter approach is used standardly when studying learning biases, e.g., when comparing length generalization scores for functions like ‘parity’, ‘majority’, ‘first’ or ‘mean’ (?, ?). Synthetic test environments have been employed also for the study of attention glitches in reasoning tasks (?). For our studies we use NT tasks, a suite of synthetic autoregressive prediction tasks which can be tuned to a desired level of difficulty, viz complexity. We will study in particular the transition between good and optimal performance when increasing context length.
3 Background
@@
Multiplying token activity with the respective query, key and value matrices generates three vectors, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$ and $\mathbf{V}_{i}$, specifically for each token $i$ in a given attention layer. The activity $\mathbf{y}_{m}$ of token $m$ is given by $\mathbf{y}_{m}=\sum_{k\leq m}a_{mk}\mathbf{V}_{k}$, when masked attention is used. The attention matrix $a_{mk}\geq 0$ is normalized row-wise, $1=\sum_{k}a_{mk}$, encoding how much information is transferred from token $k$ to token $m$. This setup implements information routing for any suitable mechanism determining the individual $a_{mk}$. The standard approach (?),
@@
takes the scalar product $\mathbf{Q}^{T}\mathbf{K}$ as the fundamental similarity measure. The softmax operation included in (1) transforms the basic similarity measure into a probability distribution function, with $Z_{m}$ being an appropriate normalization factor. One can set $\beta=1$ without affecting performance, as done here, as the standard expression $\beta=1/\sqrt{N_{\rm con}}$ only affects the scaling of entries of the key and query matrices with context length, $N_{\rm con}$ (?). For completeness we tested that the value of $\beta$ does not affect results.
@@
where $N_{m}$ is a normalization factor. An illustration is presented in Fig. 1. Importantly, EA is a function of the squared dot-product $z_{mk}$. In leading order, expressive attention is quartic in token activities and hence qualitative different from linear attention, which is bilinear in token activities.
@@
On a general level, attention mechanisms based on comparing keys and query vectors do allocate semantic meanings to the respective geometric configurations. Classically, as defined by (1), parallel/anti-parallel keys and queries correspond to high/low attention weights. Attention weights are of intermediate size for orthogonal arrangements.
Full attention matrices tend to be sparse (?), which allows in many instances to reduce the dimensionality of attention matrices from the start via a dedicated decimation process (?, ?). This observation suggest that it may be advantageous if orthogonal configurations would correspond to reduced attention weights. Eq. (2) encodes this concept, as illustrated in Fig. 1. The naming of our approach, ‘expressive attention’, is motivated by above geometric considerations.
@@ 4 Experimental settings
A basic non-linear autoregressive task is to predict the next token of a time series (?). For our investigation we are interested in a suite of prediction tasks which allows to systematically increase and tune difficulty. As a motivation we first consider the basic XOR setting (?).
@@
As a testbed, we use a standard decoder-only transformer architecture with $N_{\rm con}$ context token of embedding dimension $d=N$, where $N$ is the basis of the NT task, see Eq. (4). A straightforward orthogonal token embedding of the $N$ symbols $\{0,1,2,\dots,N\!-\!1\}$ is implemented. Attention is causal. Expressive attention works perfectly with standard positional embedding schemes, which we did however not include. This simplifies the analysis of performance as a function of input length. A single transformer bilayer is used throughout this study, with one attention head per token. Skip connections are present, with layer normalization being performed on entry, separately for the attention and the token-wise, tanh feedforward layer. The width of the feed-forward hidden layer is expanded by a factor four.
@@
Training is performed as a function of epochs, with each epoch consisting of $N_{\rm batch}$ predictions. At the start of each epoch a new random NT series is generated and encoded. The first $N_{\rm con}$ symbols are then loaded into context space. For the underlying sequence we define a batch as the task to predict one-by-one the next $N_{\rm batch}=40$ symbols. A basic SGD optimizer is used during training, with momentum $\mu=0.8$ and a learning rate $\epsilon=0.02$. Our aim is for a testbed that allows to study relative performance, in particular as a function of the complexity of the task, and not to attain optimal performance by fine-tuning meta parameters. If not stated otherwise, results shown are averaged over $16$ independent runs,
@@
Figure 3: For a NT task with basis $N=16$ and delay $\tau=2$, the performance (lines) and the loss (rescaled by a factor 0.75, circles). Result are for the standard dot-product attention (DPA, maroon), and for expressive attention (EA, blue). Performance is $1/16=0.0625$ (baseline) when predictions are random, and $100\%$ when no errors are made. Shown are results for a context length of 32 token (left panel) and 52/56 token (right panel), with DPA needing 56 token to achieve 100% performance. In this case, EA/DPA shown nearly identical training behavior.
5 Results
@@
    As a function of system size, the DPA performance rapidly increases from about $55\%$ to $100\%$. This happens between $N_{\rm con}=52$ and $56$, but it may be questionable to interpret this performance jump as an ‘emergence phenomenon’. In fact, what happens is that the structural local trap disappears when enlarging training space successively, an expected behavior.
@@
    For $\tau=2$, only $409616^{3}=$ 4096 distinct sequences exists for $N=16$. The task is hence of modest complexity, given that during training $N_{\rm batch}=40$ shifted sequences are seen for any single epoch. It is hence not surprising that the problem can be solved within 100 epochs for both models when the context length is large enough. When reducing the number adjustable parameters, both models require however larger training compute, if not failing completely. E.g., for $N_{\rm con}=16$ (not included in Fig. 3), expressive attention still converges to 100% performance, needing however about 2000 epochs.
@@
It has been pointed out that an important aspect for our understanding of transformers is the study of their loss landscape (?, ?). In this context we further investigate the structural local minima observed in Fig. 3 by comparing the two attention mechanisms for the smallest possible embedding dimension, $d=N$, which is realized for the binary case $N=2$, compare (4). An additional question here is whether EA is advantageous even in this limit.
@@
For DPA one observes extended training spans during which performance fluctuates around the baseline, 50%, essentially without improving. For EA, the equivalent training stage is substantially shorter. The phenomenology observed indicates that learning is dominated by a stochastic escape process (?). Stochastic escape is present when the additive effect of random fluctuating events allows a system to escape a local loss minimum. However, further studies would be needed to substantiate this hypothesis. Of relevance for our study is the observations that expressive attention may lead to a substantially faster convergence than the classical dot-product mechanism.
@@
[…]
@@
NT-S tasks typically lead to shorter cyclic patterns than the corresponding NT task, which makes them easier to learn. For example, the mean cycle length is 47.6 for N16T2, but only 23.8 for N16T2-S. Both systems, EA and DPA have consequently no problem to achieve 100% accuracy on N16T2-S as a single tasks. When combined with N16T2, both systems achieve still above 99% accuracy for N16T2-S, this time as subtask, as shown in Fig. 6. Also of interest is the reduction of the steady-state N16T2 performance in the mixed-task scenario, from about 0.55 to 0.34, which is also the value achieved by DPA.
In our setup, expressive attention is able to solve both tasks to about 99%, Given that N16T2 is shown only 50% of the time, it is not surprising that learning is prolonged, as evident from Fig. 6. As an experiment we reduced learning speed after 2500 epochs, finding that N16T2 performance increases a bit, by about 0.5%. Interference between competing tasks is reduced at lower learning rates.
Tasks are acquired consecutively in order of difficulty, as shown in Fig. 6, which can be interpreted as an instances of unsupervised curriculum learning (?). Our results can also be seen as an explicit example of successive learning of task quanta, as postulated in the quantization theory of neural scaling (?).
@@
[…]
@@
The data presented in Fig. 7 demonstrate that it is difficult also for expressive attention to isolate a lone if-statement, as defined by (9). Training progress is slow, achieving in the end however an average accuracy of 98.8/99.9%, respectively for $N_{\rm con}=64/128$. As usual, averages over 16 random initial conditions have been taken. The majority of runs achieves however perfect performance. We did not investigate the cause of the non-monotonic events showing up during training, which may be due to periods with an increased clustering of rare events.
@@
6 Discussion
Obtaining encouraging results, we presented a first evaluation of expressive attention (EA), focusing on single-layer models with a single attention head. We also investigate models with two or more layers and/or attention heads, together with different learning speeds and/or momentum. The results presented here are found to be robust. Additional investigation would however be necessary for a full assessment of the potential of EA, e.g. in the realm of natural language processing (?, ?), or within the context of formal languages (?, ?). Generally, we expect that expressive attention will fare at least as well as dot-product attention (DPA), a presumption that is based on the respective design principles. For DPA, large and small attention weights are constrained to a one-dimensional manifold within the space of attention heads, which is not the case for EA. However, the size of the performance boost obtainable when substituting DPA by EA may depend strongly on the use case.

### v3

@@ Abstract
Attention involves comparing query and key vectors in terms of a scalar product, $\mathbf{Q}^{T}\mathbf{K}$, together with a subsequent softmax normalization. Classically, parallel/antiparallel queries and keys lead to large/small attention weights. Here we study expressive attention (EA), which is based on $(\mathbf{Q}^{T}\mathbf{K})^{2}$, the squared dot product. In this case attention is enhanced when query and key are either parallel or antiparallel, and suppressed for orthogonal configurations. EA can be introduced into any attention-based code without additional compute costs or memory requirements. For a series of autoregressive prediction tasks, we find that EA performs at least as well as the standard mechanism, dot-product attention (DPA). Increasing task complexity, EA is observed to outperform DPA with increasing margins, which also holds for multi-task settings. For a given model size, EA manages to achieve 100% performance for a range of complexity levels not accessible to DPA.
Introduction
@@
Since its inception (Bahdanau, Cho, and Bengio 2014; Vaswani et al. 2017), the attention mechanism has had a transformative impact on machine learning (Soydaner 2022; Minaee et al. 2024). At its core, attention facilitates pairwise information routing between tokens, with information being transmitted when a given matching condition is fulfilled. For the latter query and key vectors $\mathbf{Q}$ and $\mathbf{K}$ are compared in terms of the respective dot products, $\mathbf{Q}^{T}\mathbf{K}$, which are mapped to a normalized probability distribution via a subsequent softmax operation. As illustrated in Fig. 1, this setup, dot-product attention (DPA), constrains the central part of the matching condition to a one-dimensional subspace of the otherwise potentially large space of attention heads. Given that attention matrices tend to be sparse (Likhosherstov, Choromanski, and Weller 2023), it might be favorable to allow the system to express low attention states in a larger subspace.
@@
Here we introduce and discuss a modified attention mechanism, expressive attention (EA), which is based on the presumption that low attention should correspond to orthogonal query and key configurations. This venue allows attention to express itself in the entire attention space, as illustrated in Fig. 1. All one needs to realize EA is to base attention on $(\mathbf{Q}^{T}\mathbf{K})^{2}$, the squared scalar product. In practice, this corresponds to changing a single line of code, which makes it straightforward to introduce effective attention into production codes. Computational complexity and memory requirements are not affected.
@@
We present a comparative evaluation of classical DPA and EA. For this purpose we employ a suite of autoregressive prediction tasks (NT prediction tasks) that are based on time series generated by the delayed addition of two or more numbers, where addition is modulo a generic basis $N$, like 2 or 16 (binary/hexadecimal case). NT tasks are generalized delayed XOR tasks that become increasingly difficult to solve for larger $N$ and delays $\tau$ (Schubert and Gros 2021). Models are expected to achieve not just a good performance, but 100%, which reflects the view that systems can be said to exhibit reasoning errors when achieving less than 100% accuracy for solvable tasks (Liu et al. 2024). The question is then, how long perfect performance can be maintained when systematically increasing task difficulty, viz complexity.
@@
Alternatives to the standard, attention-based transformer have been studied intensively during the past years. At a basic level, non-transformer based architectures for sequence processing have been proposed, like structured and/or selective state sequence models (Gu, Goel, and Ré 2021; Gu and Dao 2023). Other approaches aim to reduce compute costs, for example by making use of a multi-to-multi recurrent network analogy (Feng et al. 2024). Another route is linearized attention (Katharopoulos et al. 2020; Wang et al. 2020; Wu et al. 2022) for which compute scales linearly with context length, and not quadratically. In practice, linear attention comes with various problems of its own, like unbounded gradients and attention dilution (Qin et al. 2022a; Han et al. 2023). In general, linearized attention models substitute the softmax operation of classical dot-product attention by various linear kernel functions, which may have stochastic components (Choromanski et al. 2020; Peng et al. 2021), well designed spatial dependencies (Qin et al. 2022b), or being derived from a singular value decomposition (Chen et al. 2024).
@@
A different aspect is addressed by rotary positional embedding (RoPE), which adds positional information directly to the attention mechanism, via position-dependent rotations of query and key vectors (Su et al. 2021, 2024). RoPE can be used in particular for extending context windows (Chen et al. 2023). Somewhat similar in spirit is $(IA)^{3}$, which uses element-wise rescaled key and values vectors (in addition to rescaled hidden layer activities) for fine-tuning pretrained LLMs for downstream tasks (Liu et al. 2022).
@@
Two main routes are available for testing the performance of sequence processing architectures, such as transformers. The first is to use databases relevant for real-world applications (Kaddour et al. 2023; Yang et al. 2024), the second is to rely on synthetic test suites. The latter approach is used standardly when studying learning biases, e.g., when comparing length generalization scores for functions like ‘parity’, ‘majority’, ‘first’ or ‘mean’ (Abbe et al. 2023; Hahn and Rofin 2024). Synthetic test environments have been employed also for the study of attention glitches in reasoning tasks (Liu et al. 2024). For our studies we use NT tasks, a suite of synthetic autoregressive prediction tasks which can be tuned to a desired level of difficulty, viz complexity. We will study in particular the transition between good and optimal performance when increasing context length.
Background
@@
Multiplying token activity with the respective query, key and value matrices generates three vectors, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$ and $\mathbf{V}_{i}$, specifically for each token $i$ in a given attention layer. The activity $\mathbf{y}_{m}$ of token $m$ is given by $\mathbf{y}_{m}=\sum_{k\leq m}a_{mk}\mathbf{V}_{k}$, when masked attention is used. The attention matrix $a_{mk}\geq 0$ is normalized row-wise, $1=\sum_{k}a_{mk}$, encoding how much information is transferred from token $k$ to token $m$. This setup implements information routing for any suitable mechanism determining the individual $a_{mk}$. The standard approach (Vaswani et al. 2017),
@@
takes the scalar product $\mathbf{Q}^{T}\mathbf{K}$ as the fundamental similarity measure. The softmax operation included in (1) transforms the basic similarity measure into a probability distribution function, with $Z_{m}$ being an appropriate normalization factor. One can set $\beta=1$ without affecting performance, as done here, as the standard expression $\beta=1/\sqrt{N_{\rm con}}$ only affects the scaling of weight and query matrices with context length, $N_{\rm con}$ (Vaswani et al. 2017).
@@
where $N_{m}$ is a normalization factor. An illustration is presented in Fig. 1. Importantly, EA is a function of the squared dot-product $z_{mk}$. The difference between EA and DPA can be expressed in terms of attention space geometry.
@@
Attention mechanisms based on comparing keys and query vectors allocate semantic meanings to the respective geometric configurations. Classically (1), parallel/anti-parallel keys and queries lead to high/low attention weights. For a $d$-dimensional attention-head space, the d1d-1 possible orthogonal arrangements lead to intermediate attention weights.
Full attention matrices tend to be sparse (Likhosherstov, Choromanski, and Weller 2023), which allows in many instances to reduce the dimensionality of attention matrices from the start via a dedicated decimation process (Roy et al. 2021; Lou et al. 2024). This observation suggest that it may be advantageous if all possible d1d-1 dimensional orthogonal configurations would lead to reduced attention weights. Eq. (2) encodes this concept, as illustrated in Fig. 1. The naming of our approach, ‘expressive attention’, is motivated by above geometric considerations.
@@
A basic non-linear autoregressive task is to predict the next token of a time series (Li et al. 2024a). For our investigation we are interested in a suite of time series prediction tasks which allows to systematically increase and tune difficulty. As a motivation we first consider the basic XOR setting (Gros 2024).
@@
As a testbed, we use a decoder-only transformer architecture with position-specific encoding matrices, a variant denoted ‘cisformer’ (Gros 2025). We work with $N_{\rm con}$ context token and an embedding dimension $d=N$, where $N$ is the basis of the NT task, see Eq. (4). A straightforward orthogonal token embedding of the $N$ symbols $\{0,1,2,\dots,N\!-\!1\}$ is implemented. Attention is causal. Expressive attention works perfectly with standard positional embedding schemes, which we did however not include. This simplifies the analysis of performance as a function of input length. A single transformer bilayer is used throughout this study, with one attention head per token. Skip connections are present, with layer normalization being performed on entry, separately for the attention and the token-wise, tanh feedforward layer. The width of the feed-forward hidden layer is expanded by a factor four.
@@
Training is performed as a function of epochs, with each epoch consisting of $N_{\rm batch}$ predictions. At the start of each epoch a new random NT series is generated and encoded. The first $N_{\rm con}$ symbols are then loaded into context space. For the underlying sequence we define a batch as the task to predict one-by-one the next $N_{\rm batch}=40$ symbols. A basic SGD optimizer is used during training, with momentum $\mu=0.8$ and a learning rate $\epsilon=0.02$. Our aim is for a testbed that allows to study relative performance, in particular as a function of the complexity of the task, and not to attain optimal performance by fine-tuning meta parameters. If not stated otherwise, results shown are averaged over $N_{r}=16$ independent runs,
@@
Figure 3: For a NT task with basis $N=16$ and delay $\tau=2$, the performance (lines) and the loss (rescaled by a factor 0.75, circles). Result are for the standard dot-product attention (DPA, maroon), and for expressive attention (EA, blue). Performance is $1/16=0.0625$ (baseline) when predictions are random, and $100\%$ when no errors are made. Shown are results for a context length of 32 token (top panel) and 52/56 token (bottom panel), with DPA needing 56 token to achieve 100% performance. In this case, EA/DPA shown nearly identical training behavior.
Results
@@
    As a function of system size, the DPA performance rapidly increases from about $55\%$ to $100\%$. This happens between $N_{\rm con}=52$ and $56$, but it would be wrong to interpret this performance jump as an ‘emergence phenomenon’. In fact, what happens is that the structural local trap disappears eventually when enlarging training space successively, an expected behavior.
@@
    For $\tau=2$, only $16^{3}=4096$ distinct sequences exists for $N=16$. The task is hence of modest complexity, given that during training $N_{\rm batch}=40$ shifted sequences are seen for any single epoch. It is hence not surprising that the problem can be solved within 100 epochs for both models when the context length is large enough. When reducing the number adjustable parameters, both models require however larger training compute, when not failing completely. E.g., for $N_{\rm con}=16$ (not included in Fig. 3), expressive attention still converges to 100% performance, needing however about 2000 epochs.
@@
It has been pointed out that an important aspect for our understanding of transformers is the study of their loss landscape (Hahn and Rofin 2024; Choromanska et al. 2015). In this context we further investigate the structural local minima observed in Fig. 3 by comparing the two attention mechanisms for the smallest possible embedding dimension, $d=N$, which is realized for the binary case $N=2$, compare (4). An additional question is here whether EA is advantageous even in this limit.
@@
For DPA one observes extended training spans during which performance fluctuates around the baseline, 50%, essentially without improving. For EA, the equivalent training stage is substantially shorter. The phenomenology observed indicates that learning is dominated by a stochastic escape process (Gros 2024). Stochastic escape is present when the additive effect of random fluctuating events allows systems to escape a local loss minimum. However, further studies would be needed to substantiate this hypothesis. Of relevance for our studies here is the observations that expressive attention may lead to a substantially faster convergence than the classical dot-product mechanism.
@@
$\hbox{\sc NT-S:}\quad x(t)=\left(\sum_{\Delta T=1}^{\tau+1}x(t-\Delta t)\right)\%N\,$, (8)
@@
NT-S tasks typically lead to shorter cyclic patterns than the corresponding NT task, which makes them easier to learn. For example, the mean cycle length is 47.6 for N16T2, but only 23.8 for N16T2-S. Both systems, EA and DPA have consequently no problem to achieve 100% accuracy on N16T2-S as a single tasks. When combined with N16T2, both systems achieve still above 99% accuracy, as shown in Fig. 6. Also of interest is the reduction of the steady-state N16T2 performance in the mixed-task scenario, from about 0.55 to 0.34, which is also the value achieved by DPA.
In our setup, expressive attention is able to solve both tasks to about 99%, Given that N16T2 is shown only 50% of the time, it is not surprising that learning is prolonged, as evident from Fig. 6. As an experiment we lowered learning speed after 2500 epochs, finding that N16T2 performance increases a bit, by about 0.5%. Interference between competing tasks is reduced at lower learning rates.
Tasks are acquired consecutively in order of difficulty, as shown in Fig. 6, which can be interpreted as an instances of unsupervised curriculum learning (Bengio et al. 2009). Our results can be seen also as an explicit example of successive learning of task quanta, as postulated in the quantization theory of neural scaling (Michaud et al. 2024; Neumann and Gros 2022, 2024).
@@
$\hbox{\sc NT-R:}\quad\quad\left\{\begin{array}[]{ll}\mathrm{NT\!-\!S}&\mathrm{if}\ \ \ x(t-\tau-1)=0\\[2.15277pt] \mathrm{NT}&\mathrm{otherwise}\end{array}\right$. (9)
@@
The data presented in Fig. 7 demonstrate that it is difficult also for expressive attention to isolate a lone if-statement, as defined by (9). Training progress is slow, achieving in the end however an average accuracy of 98.8/99.9%, respectively for $N_{\rm con}=64/128$. As usual, averages over 16 random initial conditions have been are taken. The majority of runs achieves however perfect performance. We did not investigate the cause of the non-monotonic events showing up during training, which may be due to periods with an increased clustering of rare events.
@@
Discussion
Obtaining encouraging results, we presented a first evaluation of expressive attention (EA), focusing on single-layer models with a single attention head. We also investigate models with two or more layers and/or attention heads, together with different learning speeds and/or momentum. All results are found to be robust. Additional investigation would however necessary for a full assessment of the potential of EA, e.g. in the realm of natural language processing (Kaddour et al. 2023; Yang et al. 2024). or within the context of formal languages (Strobl et al. 2024; Abbe, Adsera, and Misiakiewicz 2023). Generally, we expect that expressive attention will fare at least as well as dot-product attention (DPA), a presumption that is based on the respective design principles. For DPA, large and small attention weights are constrained to a one-dimensional manifold within the space of attention heads, which is not the case for EA. The size of the performance boost obtainable when substituting DPA by EA may depend however strongly on the use case.